OpenStudy (anonymous):

Simplify: (24sqrt(-6)-38)/(24sqrt(-6)-108)

OpenStudy (anonymous):

\[\frac{ 24\sqrt{-6}-38 }{ 24\sqrt{-6}-108 }\]

OpenStudy (decentnabeel):

\[\frac{\sqrt{-6}\cdot \:24-38}{\sqrt{-6}\cdot \:24-108}=\frac{\sqrt{6}\cdot \:12i-19}{6\left(2\cdot \sqrt{6}i-9\right)}\]

OpenStudy (anonymous):

The first thing I would do would be to divide both numerator and denominator by 24.